HUBBLE SPACE TELESCOPE DAILY REPORT #1230 PERIOD COVERED: 0000Z Daily Status Report as of 266/0000Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5572 (WF/PC-2 Photometric Filter Calibration) The WF/PC-2 was used to make images of NGC-5139 and GRW+70D582 to make photometric calibrations for all of the WF/PC-2 filters. The observations were completed as planned, and no anomalies were reported. 1.2 Completed GHRS/2 5172 (Lyman Alpha Clouds) The GHRS/2 was used to obtain UV Spectra of a pair of quasars in Q0107-025B, separated by little more than an arcsecond. The observations were performed nominally with no reported errors. 1.3 Completed FOS/BL 5346 (B-Supergiants in Post MS Gap) The FOS (Blue Side) was utilized to look at the predicted gap in the H-R diagram of SK-69D228 to see if it is filled with early B supergiants. The instrument will obtain high S/N spectral data of four early B supergiant in the LMC to discriminate between various evolutionary scenarios and to constrain new stellar evolution scenarios. There were no reported operational errors. 1.4 Completed WF/PC-2 5655 (WF/PC-2 Flats, Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled :20 Successful :20 2.3 Significant HSTARs: None 2.4 Operations Notes: There was no significant anomalous activity in any HST onboard subsystem. A DF-224 flight software patch to change the off-nominal roll safemode limit from 35 to 40 degrees was uplinked yesterday afternoon. This was done to be consistent with the permitted operational off-nominal roll incidence envelope and the SA positional uncertainties of +/- 3.5 degrees. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /CAW -END-